OpenStudy (anonymous):

Find the unit normal vector to the surface of the given point.(Normalize the gradient vector delta F(x,y,z)) z=[squarert(x^2+y^2)] (3,4,5)

OpenStudy (anonymous):

I know, to get the level curve, I have to set the equation equal to zero so would it be: 0=[squarert(x^2+y^2-z)]? or do I just get rid of the square root altogether? 0=x^2+y^2-z^2

OpenStudy (anonymous):

I know the gradient is, delta F(x,y)=>x/(x^2+y^2)^1/2,y/(x^2+y^2)^1/2>

OpenStudy (anonymous):

trying to job my memory \[z=\sqrt{x^2+y^2}\] \[f(x,y,z)=z-\sqrt{x^2+y^2}\] want \[f(x,y,z)=0\]

OpenStudy (anonymous):

Would you mind posting your code up some other way? For some reason I can't see it, it shows this message in red text "Math Processing Error". The only way I can see it is if I refresh my page (but this is only for a couple of seconds).